OpenStudy (anonymous):

which function is odd? Check all that apply. a) y=sin(x) b) y=csc(x) c) y=sec(x) d) y=cot(x)

OpenStudy (anonymous):

i know a and b are correct. Idk about c and d

OpenStudy (anonymous):

help

OpenStudy (anonymous):

plz

OpenStudy (amoodarya):

do you know cos(x) is even or odd ?

OpenStudy (amoodarya):

so you can write sec(x)=1/cos(x)

OpenStudy (anonymous):

cos(x) is even.

OpenStudy (amoodarya):

\[f(x)=\frac{1}{cosx}\\f(-x)=\frac{1}{\cos(-x)}=\frac{1}{cosx}\] so ?

OpenStudy (anonymous):

@Agl202

OpenStudy (amoodarya):

\[cotan(x)=\frac{sinx}{cosx}\\f(-x)=\frac{\sin(-x)}{\cos(-x)}=\frac{-sinx}{cosx}=-\frac{sinx}{cosx}=-f(x)\]

OpenStudy (agl202):

y = sec x = 1/cos x substitute -x 1/ cos -x =1/cos x = sec x = y -----> even y = sin x substitute -x sin -x = - sin x = -y --------> odd y= cot x = cos x/sin x substitute -x cos -x / sin -x = cos x / (-sin x) = - cot x = -y ---------> odd y = csc x = 1 / sin x substitute -x 1/ sin -x = 1/ (-sin x)=-csc x =-y ----------> odd

OpenStudy (anonymous):

DANG.

OpenStudy (anonymous):

so the only function that is even is cos(x)?

OpenStudy (agl202):

Ya, So the other three are ur odd ones!

OpenStudy (anonymous):

so the answer is A,B,C,D?

OpenStudy (amoodarya):

note that :sec(x) is even !!!